OpenStudy (anonymous):

Rearrange F=(9/5)(C)+32 to show C=(5/9)(F-32) I'm learning about temperature in math.

OpenStudy (anonymous):

\[F=\frac{9}{5}C+32\] solve for \(C\) in two steps 1) subtract \(32\) from both sides to get \[F-32=\frac{9}{5}\]

OpenStudy (anonymous):

oops i mean \[F-32=\frac{9}{5}C\]

OpenStudy (anonymous):

then multiply both sides by \(\frac{5}{9}\) to get \[\frac{5}{9}\left(F-32\right)=C\]